web135, web136, web137, web138 abnormal server behaviour
Incident Report for Heart Internet
Resolved
This incident has been resolved.
Posted 11 days ago. Aug 12, 2017 - 13:20 BST
Update
Update - The restore is currently at 99.80% complete
Posted 11 days ago. Aug 12, 2017 - 08:24 BST
Update
The restore is currently at 86.16% complete
Posted 11 days ago. Aug 11, 2017 - 22:55 BST
Update
The restore is currently at 81.98% complete
Posted 11 days ago. Aug 11, 2017 - 21:31 BST
Update
The restore is currently at 77.49% complete
Posted 11 days ago. Aug 11, 2017 - 20:06 BST
Update
The restore is currently at 71.70% complete
Posted 11 days ago. Aug 11, 2017 - 17:49 BST
Update
The restore is currently at 68.90% complete.
Posted 11 days ago. Aug 11, 2017 - 15:41 BST
Update
The restore is currently at 67.40% complete.
Posted 11 days ago. Aug 11, 2017 - 14:31 BST
Update
The restore is currently at 65.89% complete.

Initially we expected the restoration process to take around 24 hours to be complete, but we are now seeing an increase in the estimated time until completion, as the sites being restored are incrementally larger in size.

Our current estimate for the restore is 1 AM tomorrow. We will update you as soon as we have more information.
Posted 11 days ago. Aug 11, 2017 - 13:36 BST
Update
The restore is currently at 64.70% complete.
Posted 12 days ago. Aug 11, 2017 - 12:45 BST
Update
The restore is currently at 61.68% complete.
Posted 12 days ago. Aug 11, 2017 - 11:21 BST
Update
The restore is currently at 58.38% complete.
Posted 12 days ago. Aug 11, 2017 - 09:44 BST
Update
The restore is currently at 54.81% complete.
Posted 12 days ago. Aug 11, 2017 - 08:14 BST
Update
The restore is currently at 51.67% complete.
Posted 12 days ago. Aug 11, 2017 - 07:01 BST
Update
The restore is currently at 47.62% complete.
Posted 12 days ago. Aug 11, 2017 - 06:01 BST
Update
The restore is currently at 45.45% complete.
Posted 12 days ago. Aug 11, 2017 - 05:03 BST
Update
The restore is currently at 43.43% complete.
Posted 12 days ago. Aug 11, 2017 - 04:08 BST
Update
The restore is currently at 41.40% complete.
Posted 12 days ago. Aug 11, 2017 - 03:01 BST
Update
The restore is currently at 39.58% complete.
Posted 12 days ago. Aug 11, 2017 - 01:45 BST
Update
The restore is currently at 37.34% complete.
Posted 12 days ago. Aug 11, 2017 - 00:49 BST
Update
The restore is currently at 34.85% complete.
Posted 12 days ago. Aug 10, 2017 - 23:24 BST
Update
The restore is currently at 34.05% complete.
Posted 12 days ago. Aug 10, 2017 - 22:35 BST
Update
The restores are currently on 31%, the ETA remains unchanged currently.
Posted 12 days ago. Aug 10, 2017 - 21:37 BST
Update
The restorations are currently at 23% complete, we will update on this progress throughout the night.
Posted 12 days ago. Aug 10, 2017 - 19:19 BST
Update
18% of all the packages on this cluster have been restored now. These sites are back online and should be fully functional.

We still expect this to finish according to the ETA provided. We'll update you as soon as we have more information.
Posted 12 days ago. Aug 10, 2017 - 18:01 BST
Update
The restoration process is 15% complete now and all sites that have been restored so far are back online. There are still no changes to the ETA provided for this restoration.

We'll update you as soon as we have more information.
Posted 12 days ago. Aug 10, 2017 - 17:06 BST
Update
The restoration is still progressing as planned on the NAS drives, and there are no changes to the ETA at the moment.

We are prioritising small sized sites at the moment in order to ensure that the majority of sites on these servers are back online as quickly as possible.

We'll update you as soon as we have more information.
Posted 12 days ago. Aug 10, 2017 - 16:07 BST
Identified
An issue has been identified with the NAS drives for the above servers. All sites on these servers are effected by this issue, but the databases remain intact.

We are restoring the NAS drives from our backups and we expect this restoration process to take around 24 hours.

We'll update you as soon as we have more information, we apologise for any inconvenience caused regarding this.
Posted 12 days ago. Aug 10, 2017 - 14:19 BST
Investigating
We are currently seeing abnormal behaviour on this server, and are investigating the causes.
We expect to have it back to normal operation soon, and apologise for any inconvenience caused.
Posted 13 days ago. Aug 10, 2017 - 11:42 BST
This incident affected: Webservers and VPS & Hybrid Servers.